An article in the ACI Materials Journal (Vol. 84, 1987, pp. 213-216) describes several experiments investigating the roddingNormal Probability Plot of the Residuals (response is STRENGTH) Residuals Versus RODDING (response 18 STRENGTH) 100- Normal SResidualsVersusthe Fitted Values (response is STRENGTH) 100 Residual -100 1 500 1550 1600 Fitted Value
Answer #1

(A)

(1) Null and Alternative Hypotheses

The following null and alternative hypotheses need to be tested:

Ho: μ1 = μ2 = μ3 = μ4 or no difference in the rodding level.

Ha: Not all means are equal or the rodding level of atleast two groups differ.

The above hypotheses will be tested using an F-ratio for a One-Way ANOVA.

(2) Rejection Region

Based on the information provided, the significance level is \alpha = 0.05α=0.05, and the degrees of freedom are df_1 = 3df1​=3 and df_2 = 3df2​=3, therefore, the rejection region for this F-test is R={F:F>Fc​=4.066}.

(3) Test Statistics

The following table is obtained:

Group 1 Group 2 Group 3 Group 4
1530 1610 1560 1500
1530 1650 1730 1490
1440 1500 1530 1510
Sum = 4500 4760 4820 4500
Average = 1500 1586.667 1606.667 1500
\sum_i X_{ij}^2 =∑i​Xij2​= 6755400 7564600 7767400 6750200
St. Dev. = 51.962 77.675 107.858 10
SS = 5400 12066.666666667 23266.666666667 200
n = 3 3 3 3

The total sample size is N = 12N=12. Therefore, the total degrees of freedom are:

dftotal​=12−1=11

Also, the between-groups degrees of freedom are dfbetween​=4−1=3, and the within-groups degrees of freedom are:

dfwithin​=dftotal​−dfbetween​=11−3=8

First, we need to compute the total sum of values and the grand mean. The following is obtained

∑​Xij​=4500+4760+4820+4500=18580

Also, the sum of squared values is

∑​Xij2​=6755400+7564600+7767400+6750200=28837600

Based on the above calculations, the total sum of squares is computed as follows

1 185802 SSIotal = ΣΧ ΣΧ, = 28837600 - 69566.667 N 12 1, 2.]

The within sum of squares is computed as shown in the calculation below:

SS within 200 = 40933.333 S Swithin groups 5400+12066.666666667+23266.666666667+

S Sbetween 28633.333 M Sbetween - 9544.444 dfbetween 3

40933.333 MS within S Swithin dfurithin 5116.667 8

F M Sbetween M Swithin 9544.444 5116.667 1.865

F(critical)=4.066

(4) Decision about the null hypothesis

Since it is observed that F=1.865≤Fc​=4.066, it is then concluded that the null hypothesis is not rejected.

Using the P-value approach: The p-value is p=0.2138, and since p=0.2138≥0.05, it is concluded that the null hypothesis is not rejected.

(5) Conclusion

It is concluded that the null hypothesis Ho is not rejected. Therefore, there is not enough evidence to claim that not all 4 population means are equal, at the α=0.05 significance level or or no difference in the rodding level

(B) P-value=0.2138

.

i hope it helps..

If you have any doubts please comment and please don't dislike.

PLEASE GIVE ME A LIKE. ITS VERY IMPORTANT FOR ME.